Project

General

Profile

NDOS Commissioning Meeting 2011/08/22

Attendance: Jon, Susan, Gary, Leon, Sue, Jarek, Minerba, Martin, Enrique, Rafael, Mark Mathis, Mark Messier, Gavin, Tim

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • From Mat:

"I'll be traveling during the meeting this morning. The only things I wanted to bring up were that we did enable cooling on 19 good shimmed APDs. We should take a DSO scan this morning to see how they're doing but the preliminary report from the shifters based on rate in the OnMon is that 3 of those 19 may have gone bad over the weekend.

Also on Friday the shifter reported that one of the golden APD had gone into alarm on the cooling GUI. We found that this unit was drawing maximal TEC current but only cooling to -5 C. The FEB temp was also very high. Martin can say a bit more about the history of this based on what we see logged in the database. I went to the hall to visually inspect (Unit 3-2-2 19) and found that many of the conductors on the black TEC lead wire were broken at the connector. It looks like this unit first started acting up just after ferrite beads were installed so that is most likely when the damage occurred. Denis can repair this easily once we return. The APD on this channel still looked good in the most recent DSO scan. For now we disabled cooling on that channel. It would be nice to have alarms on the drive current and the FEB temp as soon as possible, those stats would have revealed this issue well before the APD temperature went funny."

  • Martin will work on implementing alarms on TEC drive current and FEB temperature. Should be done by the end of the week.

DAQ Status & Plans

  • Leon:
    • 13 APDs have gone through "soak" tests (overnight tests) and look very good. Many others were tested; about ~6 look to be "B" grade.
    • vacuum test for seal between carrier board and spacer frame is being developed.
    • expects to ship back 13 "A" grade by end of week
    • recent failed golden APDs (with heatsink attached) should arrive today.
    • unused master was removed from NDOS and put on test-stand, test-stand TDU is being shipped up to Ash River.
  • Peter: test-stand MTDU has more stable oscillator and so is being shipped up Ash River
  • Software:
    • Peter: Ron has a new kernel module that should fix the external timing problem on the DCMs
    • Sue: dcmapplication is not able to run in external timing for more than a few hundred microslices on the test-stand. Did the TDU test-stand version change? Peter: quite possibly. We will roll back naturally since the TDU from NDOS is being put in place at the test-stand. Sue will look to see if there are instructions somewhere on how to roll back the firmware.
  • Peter: Neal already moved TDU and updated the firmware. Sue will follow up with him.

DCS Status & Plans

  • Aside from working on the TEC drive current and FEB temp. alarms:
  • Martin:
    • new power-on/off and hv-on/off scripts to distinguish between "all" (for real) vs. "
    • pedestalrunner still uses the scripts in usr/local/bin
    • would like to test IOCs to control power supplies on teststand. Will email nova-daq group when ready to test.
    • Zukai would like to test GUI layout of FEBs for temp. monitoring. Would like to test on novatest01. Jon: should be able to test GUI on one of the DCS machines.

AOB

  • None